A cube of 2.13 inches on the side has a mass of 36.6 g. What is the density in ( g)/(c)m3? 1 inch =2.54 cm


1
Expert's answer
2021-01-07T05:32:15-0500

Side = 2.13 inch

1 inch = 2.54 cm

Side = 2.13 inch = 2.13×2.54 = 5.4102 cm

mass = 36.6g

Volume = side3 = (5.4102)3 = 158.36 cm3

Density = "\frac{mass}{volume}" = "\frac{36.6g}{158.36cm^3}"


Density = 0.2311 g/cm3